# MathDuels: Evaluating LLMs as Problem Posers and Solvers

### 23 Apr 2026 v1

@@ 1 Introduction
Concretely, a single evaluation round proceeds as follows. Each of $N$ participating models authors $K$ problems via a prompted generation pipeline that includes adversarial difficulty amplification. Every non-author model then attempts each problem. Answers are judged automatically through symbolic verification, and problems that defeat any solver undergo a validity check to filter ill-posed or ambiguous questions. The resulting outcome matrix feeds a Rasch model that jointly estimates solver abilities and problem difficulties, with author quality derived from the difficulties of each model’s authored problems.
@@ 2 The MathDuels protocol
We formalize an analogous protocol for language models. A MathDuels run takes a set of $N$ models and subjects each to evaluation in two roles: as an author of math problems and as a solver of problems authored by others. The protocol proceeds in four phases (generation, solving, judging, and ranking), each designed to be fully automated and reproducible without human annotation in the loop. Algorithm 1 gives a high-level pseudocode summary; the subsections that follow expand each phase.
@@ 2.1 Arena setup
Let $\mathcal{M}=\{m_{1},\dots,m_{N}\}$ be the set of participating models and let $K$ be the number of problems each model generates, giving $NK$ problems in total. Every model attempts every problem it did not author, so for each solver–problem pair $(m,p)$ with $m\neq\mathrm{author}(p)$ we observe a binary outcome
@@ 2.4 Verification
When at least one solver answer is incorrect, we run a secondary verification step. Let $\mathcal{A}_{p}=\{(a_{i},r_{i})\}$ denote the set of unique solver answers and their reasoning traces for problem $p$. Verification is triggered only when disagreement exists:
@@ 3.1 Setup
Each model authors 30 problems per run, uniformly distributed across six broad mathematical domains – analysis, algebra, geometry & topology, discrete mathematics, probability & statistics, and applied & computational mathematics – via domain-specific meta-prompts (Appendix C.1). Every problem is generated from the 3-step protocol introduced in Section 2.2. We use math-verify (Kydlíček, 2025) for expression equivalence checking. We select GPT-5.4-high as backbone of verifier, which is ablated in Section 4.3. All inference calls use temperature 1.0.
@@ 3.2 Results
The Rasch model (Section 2.5) jointly estimates solver abilities and problem difficulties, with Grok-4.1-fast-high anchored at 1500. Figure 2 shows the composite rating and its bootstrapped 95% confidence intervals. Interestingly, the best solver is not the best overall model. GPT-5.4-high leads all participants in solver rating, yet Gemini-3.1-Pro-high takes the top composite rank because it is by far the strongest author in the field: its authored problems have the lowest average solve rate of any participant (62.9%). This authoring advantage is invisible to any solve-only benchmark. Notably, Grok-4.20-high has the largest gap in solver and author rating. We analyzed its authored problems along with Gemini-3.1-Pro-high’s in Section 3.3.
@@ 3.3 Problem analysis
For example, GPT-5.2-high authored a topology problem on the invariant-factor decomposition of $\pi_{2}$ of a mapping torus built from $\bigvee_{i=1}^{4}S_{i}^{2}$: heavy machinery, 17/19 solve rate. The same model posed a combinatorics question asking for the number of $3$-colorings of a regular $12$-gon under dihedral symmetry with exact color-count and adjacency constraints: two sentences to state, 11/19 solve rate. The topology problem advertises its difficulty; the combinatorics problem hides it (see Appendix C.1 for further analysis).
@@ C.1 Problem Taxonomy
be the wedge of four copies of the $2$-sphere, form a simply connected CW-complex $X$ by attaching a $3$-cell along the class $2\alpha_{1}+2\alpha_{3}\in\pi_{2}(F,*)$, and let $f:(X,*)\to(X,*)$ be a based cellular homotopy equivalence inducing a specified homomorphism on $\pi_{2}(X,*)$. Define the mapping torus
@@
“Consider a regular $12$-gon with vertices colored red, blue, and green so that exactly four vertices have each color, adjacent vertices are never the same color, and for each ordered pair of distinct colors $(X,Y)$, the number of indices $i$ such that vertex $i$ has color $X$ and vertex $i+1$ has color $Y$ is exactly $2$. Two colorings are identified up to dihedral symmetry. How many distinct colorings are there?”
@@
“A standard 52-card deck is shuffled into a uniformly random order. Define event $A$ as the event that the first King in the deck appears in an earlier position than the second Queen (i.e., the Queen whose position is second-smallest among all four Queens), and define event $B$ as the event that the first King in the deck appears in an earlier position than the first Jack. Compute $P(A\mid B)$.”

### 27 Apr 2026 v2

@@ 1 Introduction
Concretely, a single evaluation round proceeds as follows. Each of NN participating models authors $K$ problems via a prompted generation pipeline that includes adversarial difficulty amplification. Every non-author model then attempts each problem. Answers are judged automatically through symbolic verification, and problems that defeat any solver undergo a validity check to filter ill-posed or ambiguous questions. The resulting outcome matrix feeds a Rasch model that jointly estimates solver abilities and problem difficulties, with author quality derived from the difficulties of each model’s authored problems.
@@ 2 The MathDuels protocol
We formalize an analogous protocol for language models. A MathDuels run takes a set of NN models and subjects each to evaluation in two roles: as an author of math problems and as a solver of problems authored by others. The protocol proceeds in four phases (generation, solving, judging, and ranking), each designed to be fully automated and reproducible without human annotation in the loop. Algorithm 1 gives a high-level pseudocode summary; the subsections that follow expand each phase.
@@ 2.1 Arena setup
Let $\mathcal{M}=\{m_{1},\dots,m_{N}\}$ be the set of participating models and let $K$ be the number of problems each model generates, giving NKNK problems in total. Every model attempts every problem it did not author, so for each solver–problem pair $(m,p)$ with $m\neq\mathrm{author}(p)$ we observe a binary outcome
@@ 2.4 Verification
When at least one solver answer is incorrect, we run a secondary verification step. Let $\mathcal{A}_{p}=\{(a_{i},r_{i})\}$ denote the set of unique solver answers and their reasoning traces for problem pp. Verification is triggered only when disagreement exists:
@@ 3.1 Setup
Each model authors 30 problems per run, uniformly distributed across six broad mathematical domains – analysis, algebra, geometry & topology, discrete mathematics, probability & statistics, and applied & computational mathematics – via domain-specific meta-prompts (Appendix C.1). Every problem is generated from the 3-step protocol introduced in Section $2.$ We use math-verify (Kydlíček, 2025) for expression equivalence checking. We select GPT-5.4-high as backbone of verifier, which is ablated in Section 4.3. All inference calls use temperature 1.0.
@@ 3.2 Results
The Rasch model (Section 2.5) jointly estimates solver abilities and problem difficulties, with Grok-4.1-fast-high anchored at 1500. Figure 2 shows the composite rating and its bootstrapped 95% confidence intervals. Interestingly, the best solver is not the best overall model. GPT-5.4-high leads all participants in solver rating, yet Gemini-3.1-Pro-high takes the top composite rank because it is by far the strongest author in the field: its authored problems have the lowest average solve rate of any participant (62.9%). This authoring advantage is invisible to any solve-only benchmark. Notably, Grok-4.20-high has the largest gap in solver and author rating. We analyzed its authored problems along with Gemini-3.1-Pro-high’s in Section $3.$
@@ 3.3 Problem analysis
For example, GPT-5.2-high authored a topology problem on the invariant-factor decomposition of $\pi_{2}$ of a mapping torus built from $\bigvee_{i=1}^{4}S_{i}^{2}$: heavy machinery, 17/19 solve rate. The same model posed a combinatorics question asking for the number of 33-colorings of a regular $12$-gon under dihedral symmetry with exact color-count and adjacency constraints: two sentences to state, 11/19 solve rate. The topology problem advertises its difficulty; the combinatorics problem hides it (see Appendix C.1 for further analysis).
@@ C.1 Problem Taxonomy
be the wedge of four copies of the 22-sphere, form a simply connected CW-complex $X$ by attaching a 33-cell along the class $2\alpha_{1}+2\alpha_{3}\in\pi_{2}(F,*)$, and let $f:(X,*)\to(X,*)$ be a based cellular homotopy equivalence inducing a specified homomorphism on $\pi_{2}(X,*)$. Define the mapping torus
@@
“Consider a regular $12$-gon with vertices colored red, blue, and green so that exactly four vertices have each color, adjacent vertices are never the same color, and for each ordered pair of distinct colors $(X,Y)$, the number of indices $i$ such that vertex $i$ has color $X$ and vertex $i+1$ has color $Y$ is exactly 22. Two colorings are identified up to dihedral symmetry. How many distinct colorings are there?”
@@
“A standard 52-card deck is shuffled into a uniformly random order. Define event AA as the event that the first King in the deck appears in an earlier position than the second Queen (i.e., the Queen whose position is second-smallest among all four Queens), and define event $B$ as the event that the first King in the deck appears in an earlier position than the first Jack. Compute $P(A\mid B)$.”